# Bridging Operator Learning and Conditioned Neural Fields: A Unifying Perspective

### v1

@@ Abstract
Operator learning is an emerging area of machine learning which aims to learn mappings between infinite dimensional function spaces. Here we uncover a connection between operator learning architectures and conditioned neural fields from computer vision, providing a unified perspective for examining differences between popular operator learning models. We find that many commonly used operator learning models can be viewed as neural fields with conditioning mechanisms restricted to point-wise and/or global information. Motivated by this, we propose the Continuous Vision Transformer (CViT), a novel neural operator architecture that employs a vision transformer encoder and uses cross-attention to modulate a base field constructed with a trainable grid-based positional encoding of query coordinates. Despite its simplicity, CViT achieves state-of-the-art results across challenging benchmarks in climate modeling and fluid dynamics. Our contributions can be viewed as a first step towards adapting advanced computer vision architectures for building more flexible and accurate machine learning models in physical sciences.
00footnotetext: Our code is available at https://github.com/PredictiveIntelligenceLab/cvit
@@ 1 Introduction
Neural fields have recently gained prominence in computer vision as a memory-efficient approach for representing high-resolution images and scenes. At their core, neural fields are simply functions over a continuous domain that are parameterized by neural networks [1]. A key advance has been the development of conditioned neural fields, which allow these models to change based on auxiliary inputs. This is achieved through a base field combined with conditioning mechanisms that modulate the field’s forward pass to encode multiple signals into a single neural network architecture [2].
In parallel, the field of scientific machine learning has seen the emergence of neural operators – a class of deep learning models well-suited for operating on functional data common in physical sciences applications, such as building surrogates for expensive partial differential equation (PDE) solvers. For example, neural operators can learn mappings between infinite-dimensional function spaces to accelerate fluid dynamics simulations [3]. The formulation of widely used neural operator architectures like Fourier Neural Operators (FNO) [4, 5] and DeepONet [6, 7] is primarily motivated by mathematical intuition about the structure of the spaces that PDE solutions live in, These architectures are often supported by universal approximation guarantees [8, 9, 10, 11], ensuring their theoretical robustness. Furthermore, these models can also be interpreted as learning low-dimensional manifolds in function spaces [12, 13]. However, from a deep learning perspective, the architectural choices in current neural operator models appear quite distant to the wealth of techniques employed in other domains, potentially limiting their performance.
In this work, we uncover a previously unexplored connection between operator learning architectures and conditioned neural fields, providing a unifying framework to compare and contrast popular operator learning methods. Specifically, our main contributions can be summarized as follows:
  • Neural operators as conditioned neural fields. We observe that many widely used operator learning models can be reformulated as neural fields with conditioning mechanisms restricted to pointwise and/or global information.
  • A unifying perspective. We provide a general rubric for examining and contrasting neural operators through their respective choices of conditioning mechanisms and base fields.
  • Introducing the Continuous Vision Transformer (CViT). We put forth a novel neural operator that employs a vision transformer encoder and cross-attention to modulate a base field built with a trainable grid-based positional encoding of query coordinates.
  • Simplicity meets state-of-the-art performance. We demonstrate that, despite its simplicity, CViT achieves state-of-the-art performance on challenging benchmarks in climate modeling and fluid dynamics.
Taken together our results demonstrate the potential of adapting advanced computer vision architectures to develop more flexible and accurate machine learning models for physical sciences. Our contributions represent a rigorous initial step towards bridging the gap between the fields of computer vision and scientific machine learning, and aim to spark further research in this direction.
2 Neural Fields
A neural field is a function over a continuous domain parameterized by a neural network [1]. Recently, the computer vision community has adopted neural fields to represent 3D spatial fields, such as neural radiance fields (NeRFs) [14], spatial occupancy fields [15], and signed distance functions [16]. These fields represent functions defined over continuous spatial domains, with NeRFs additionally taking an element of a continuous ray space as input. In this paper, a point in the domain of a neural field is referred to as a query point, here denoted by $y$. Conditioned neural fields allow these fields to change based on auxiliary inputs $u$, often represented as latent codes $z$, without the need to retrain the entire field; see Figure 1. The latent code affects the forward pass by modifying the parameters of the base field through either global conditioning, where the forward pass is affected uniformly across all query points, or local conditioning, where the modifications depend on the specific query point.
Global Conditioning.
Global conditioning in a neural field refers to modifying the forward pass only by the input $u$ and in a uniform manner over the entire query domain. This can be achieved through concatenation, where a global latent code is appended to the query vector and passed together through an MLP [17, 18, 19]. Alternatively, a hypernetwork could accept the auxiliary input and determine the neural field’s parameter values based on this input [20]. However, modifying every parameter of the base field for different auxiliary inputs can be computationally expensive. For example, feature-wise Linear Modulation (FiLM) [21], which originates from conditional batch normalization [22], has been proposed to modify only the sine activation function parameters in a SIREN network [23]. This technique has been used for global conditioning in methods such as $\pi$-GAN [24].
Local Conditioning.
Local conditioning occurs when the latent code $z$, which modifies the base field’s parameters, is a function of the input query $y$, i.e., $z=g(y)$. This results in query-dependent field parameters that can vary across the query domain. A common approach to generate local latent codes is to process the image or scene with a convolutional network and extract features from activations with receptive fields containing the query point. This can be done using 2D convolutions of projected 3D scenes [25, 26, 27] or 3D convolutions [28, 29, 30]. Muller et al. [31] learn features directly on multiresolution grid vertices and create local features via nearest grid point interpolation for each grid. Chan et al. [32] create features along three 2D planes and interpolate features from projections onto these planes. DeVries et al. [33] learn latent codes over patches on a 2D “floorplan” of a scene. Other methods divide the query domain into patches or voxels and learn a latent code $z$ for each patch or voxel [34, 35, 36].
3 Operator Learning Architectures as Neural Fields
Here we demonstrate that many popular operator learning architectures are in fact neural fields with local and/or global conditioning. While connections between architectures such as the DeepONet and the Fourier Neural Operator have been described previously [10], they are typically formulated under certain architecture choices which allow these architectures to approximate one another. In contrast, here we show that, by viewing these architectures as conditioned neural fields, they are explicitly related through their respective choices of conditioning mechanisms and base fields.
3.1 Encoder-Decoder Neural Operators
A large class of operator learning architectures can be described as encoder-decoder architectures according to the following construction. To learn an operator ${\mathcal{G}}:{\mathcal{X}}\to{\mathcal{Y}}$, an input function $u\in{\mathcal{X}}$ are first mapped to a finite dimensional latent representation through the encoding map ${\mathcal{E}}:{\mathcal{X}}\to{\mathbb{R}}^{n}$. This latent representation is then mapped to a callable function through a decoding map ${\mathcal{D}}:{\mathbb{R}}^{n}\to{\mathcal{Y}}$. The composition of these two maps gives the architecture of the operator approximation scheme, ${\mathcal{G}}={\mathcal{D}}\circ{\mathcal{E}}$. The class of encoder-decoder architectures includes the DeepONet [6], the PCA-based architecture of [37], and NoMaD [12].
DeepONet.
For example, in a DeepONet, the base field is a neural network $t_{\theta}(y)$ (trunk network) appended with a last linear layer, and the decoder map conditions the weights of this last layer via the output of the encoding map ${\mathcal{E}}_{\phi}(u)$ (branch network); see Figure 2(a) for a visual illustration. In other words, the decoder map is a weighted linear combination of $n$ basis elements $t_{\theta}(y)$, where the weights are modulated by a global encoding of the input function ${\mathcal{E}}_{\phi}(u)$, yielding outputs of the form $s(y)=\langle{\mathcal{E}}_{\phi}(u),t_{\theta}(y)\rangle$. The PCA-based architecture of [37] can be viewed analogously using as basis elements the leading $n$ eigenfunctions of the empirical covariance operator computed over the output functions in the training set.
NoMaD.
Seidman et al. [12] demonstrated how encoder-decoder architectures can employ nonlinear decoders to learn low-dimensional manifolds in an ambient function space. These so-called NoMaD architectures typically perform their conditioning by concatenating the input function encoding with the query input, i.e., $s(y)=f_{\theta}(y,{\mathcal{E}}_{\phi}(u))$; see Figure 2(b) for an illustration.
3.2 Integral Kernel Neural Operators
[…]
Graph Neural Operator (GNO).
A single layer of the GNO [38] is given by
where $W\in{\mathbb{R}}^{d_{u}\times d_{s}}$ is a trainable weight matrix, and $\kappa_{\phi}(\cdot)$ is a local message passing kernel parameterized by $\phi$ [38]. As illustrated in Figure 2(c), the GNO layer is a conditioned neural field with local and global conditioning on the input function $u$. The base field uses a fixed positional encoding $\gamma(y)$ that maps query coordinates $y$ to input function values at neighboring nodes $x\in N(y)$, i.e., $\gamma(y)=u(x)/|N(y)|$, followed by a linear layer. The encoding width equals the maximum number of neighboring nodes, with zeros filling tailing entries for queries with fewer neighbors. The GNO layer is conditioned globally via the parameterized kernel $\kappa_{\phi}(\cdot)$ modulating the linear layer weights, and locally via a position-depended bias term $Wu(y)$ added as a skip connection. The layer outputs $s(y)$ are finally obtained by summing the linear layer outputs and the skip connection, then applying a nonlinear activation function $\sigma$.
Fourier Neural Operator (FNO).
The FNO introduced a computationally efficient approach for performing global message passing in GNO layers assuming a stationary kernel $\kappa_{\phi}(\cdot)$ [4, 39]. Leveraging the Fourier transform, a single layer of FNO [4] is given by
where $W\in{\mathbb{R}}^{d_{o}\times d_{o}}$ and $K\in{\mathbb{R}}^{n\times n}$ are trainable weight matrices, ${\mathcal{F}}_{n}:L^{2}\to{\mathbb{C}}^{n}$ is the Fourier transform truncated to the first $n$ modes, and ${\mathcal{F}}_{n}^{-1}:{\mathbb{C}}^{n}\to L^{2}$ is the left inverse of this operator. We show that the FNO layer can be viewed as a conditioned neural field employing both local and global conditioning on the input function $u$. Without loss of generality, we demonstrate this for a one-dimensional query domain $y\in{\mathbb{R}}$. To this end, note that the convolutional part of the FNO layer (equation 2) can be re-written as (see Appendix B for a derivation)
Extending to other neural operators.
In practice, GNO and FNO layers are stacked to form deeper architectures with a compositional structure. These can also be interpreted as conditioned neural fields, where the base field corresponds to the last GNO or FNO layer, while all previous layers are absorbed into the definition of the conditioning mechanism. Analogously, one can examine a broader collection of models that fall under the class of Integral Kernel Neural Operators, such as [40, 41, 42, 43, 44].
@@ 4 Continuous Vision Transformer (CViT)
While the mathematical foundations of neural operators are well-established [8, 9, 10, 11], the architectural design of these models has not kept pace with the rapid advancements in deep learning. As shown in the previous section, the most common neural operator architectures, such as GNO, FNO, DeepONet, and NoMaD, essentially rely on simple building blocks like Fourier layers or fully connected networks, which may not effectively capture the complex spatial dependencies and multiscale features present in many physical systems. In contrast, computer vision has seen a proliferation of sophisticated architectures, such as convolutional neural networks and vision transformers, which have been highly successful in learning rich representations from complex high-dimensional data. This suggests that there is significant potential to improve the performance and expressiveness of neural operators by drawing inspiration from these advanced techniques.
By using the perspective of neural fields as a bridge to these computer vision architectures, we introduce the Continuous Vision Transformer (CViT), a novel neural operator architecture that combines the strengths of vision transformers and neural fields to learn more powerful and flexible representations of time-dependent physical systems. The CViT can be viewed as a well-defined conditioned neural field with transformer architectures being the base field and conditioning function. As we will see in detail in the following paragraphs, the base field comprises of a trainable positional encoding of query coordinates (see equation 5), followed by a standard Transformer decoder. This base field transformer is then globally conditioned by cross-attention, using the output of a standard ViT backbone as key and values, which is evaluated on the input function $u$. See Figure 3 for a visual summary.
Base field of CViT.
The base field of CViT consists of three components: a positional embedding, a transformer backbone, and a nonlinear projection mapping. To create the positional encoding, we start by creating a uniform grid $\{\mathbf{y}_{ij}\}\subset[0,1]^{2}$, for $i=1,\dots N_{x}$ and $j=1,\dots N_{y}$, along with associated trainable latent grid features $\mathbf{x}\in{\mathbb{R}}^{N_{x}\times N_{y}\times C}$ [28, 45, 46]. Here $\mathbf{x}$ is initialized by a unit Gaussian distribution, and the hyper-parameters $N_{x},N_{y}$ are typically set to match the resolution of the target output functions. For a single arbitrary query point $y\in{\mathbb{R}}^{2}$, we then compute a Nadaraya-Watson interpolant [47, 48] over grid latent features $\mathbf{x}$ as follows,
$\mathbf{x}^{\prime}=\sum_{i=1}^{N_{x}}\sum_{j=1}^{N_{y}}w_{ij}\mathbf{x}_{ij},\quad w_{ij}=\frac{\exp\left(-\epsilon\|y-\mathbf{y}_{ij}\|^{2}\right)}{\sum_{ij}\exp\left(-\epsilon\|y-\mathbf{y}_{ij}\|^{2}\right)},$ (5)
where $\epsilon>0$ is a hyper-parameter that determines the locality of the interpolated features. The positional encoding from the interpolated grid feature $\mathbf{x}_{0}=\mathbf{x}^{\prime}\in{\mathbb{R}}^{1\times C}$ is then fed as query inputs to the transformer backbone, where we globally condition its output with a latent variable $\mathbf{z}_{L}(u)$ of the input $u$ (to be described next) as keys and values, applying $K$ cross-attention Transformer blocks as

### v2

# CViT: Continuous Vision Transformer for Operator Learning

@@ Abstract
Operator learning, which aims to approximate maps between infinite-dimensional function spaces, is an important area in scientific machine learning with applications across various physical domains. Here we introduce the Continuous Vision Transformer (CViT), a novel neural operator architecture that leverages advances in computer vision to address challenges in learning complex physical systems. CViT combines a vision transformer encoder, a novel grid-based coordinate embedding, and a query-wise cross-attention mechanism to effectively capture multi-scale dependencies. This design allows for flexible output representations and consistent evaluation at arbitrary resolutions. We demonstrate CViT’s effectiveness across a diverse range of partial differential equation (PDE) systems, including fluid dynamics, climate modeling, and reaction-diffusion processes. Our comprehensive experiments show that CViT achieves state-of-the-art performance on multiple benchmarks, often surpassing larger foundation models, even without extensive pretraining and roll-out fine-tuning. Taken together, CViT exhibits robust handling of discontinuous solutions, multi-scale features, and intricate spatio-temporal dynamics. Our contributions can be viewed as a significant step towards adapting advanced computer vision architectures for building more flexible and accurate machine learning models in the physical sciences.
sifan.wang@yale.edu{seidj,wangh19}@sas.upenn.edu{shyamss,pappasg,pgp}@seas.upenn.edu
@@ 1 Introduction
Neural operators (Kovachki et al., 2023) have emerged as a powerful class of deep learning models for scientific machine learning applications, particularly in building efficient surrogates for expensive partial differential equation (PDE) solvers. These models aim to learn mappings between infinite-dimensional function spaces, enabling rapid acceleration of complex simulations in fields such as fluid dynamics and climate modeling (Azizzadenesheli et al., 2024; Serrano et al., 2023). Popular architectures like Fourier Neural Operators (FNO) (Li et al., 2021; Tran et al., 2021) and DeepONet (Lu et al., 2021; Wang et al., 2022) have shown promising results, supported by universal approximation guarantees (Chen & Chen, 1995; Lanthaler et al., 2022; Kovachki et al., 2021; De Ryck & Mishra, 2022) and interpretations as learning low-dimensional manifolds in function spaces (Seidman et al., 2022; 2023).
However, the architectural design of current neural operators has largely been motivated by mathematical intuition about the structure of PDE solution spaces, rather than leveraging recent advances in deep learning architectures. This approach, while theoretically sound, may limit the performance and flexibility of these models when applied to complex, real-world physical systems.
In parallel, the field of computer vision has seen significant advancements in neural field models, which represent continuous functions parameterized by neural networks (Stanley, 2007). A key innovation in this domain has been the development of conditioned neural fields, which allow for auxiliary inputs to modulate the field’s behavior (Xie et al., 2022). These models have demonstrated remarkable effectiveness in tasks such as representing high-resolution images and 3D scenes (Mildenhall et al., 2021; Müller et al., 2022).
Motivated by the success of conditioned neural fields in computer vision, we introduce the Continuous Vision Transformer (CViT), a novel neural operator architecture that bridges the gap between advanced computer vision techniques and scientific machine learning. Our main contributions can be summarized as follows:
  • Continuous Vision Transformer (CViT). We introduce CViT, a neural operator that combines a vision transformer encoder with a novel, trainable grid-based coordinate embedding in its decoder. This design enables CViT to effectively capture multi-scale spatial dependencies and allows for continuous querying at arbitrary resolutions.
  • Theoretical Insights into Lipschitz Constant and Spectral Bias. We present novel insights into the impact of coordinate embeddings on the Lipschitz constant of neural fields and neural operators. This analysis supports the choice of the proposed grid-based coordinate embedding and its efficacy in mitigating spectral bias, a common limitation in training neural operators .
  • State-of-the-art Performance. Despite its architectural simplicity, CViT achieves state-of-the-art results on challenging benchmarks in climate modeling, fluid dynamics and reaction-diffusion processes. It often outperforms larger models while using fewer parameters and without requiring extensive pretraining.
Taken together, we explore the connection between operator learning architectures and conditioned neural fields, offering a unifying perspective for examining differences among popular operator learning models (see Appendices B and C). This perspective not only enhances understanding of existing architectures but also provides a general framework for designing future neural operators. By bridging the gap between computer vision and scientific machine learning, CViT paves the way for improved surrogates of complex physical systems, with potential applications ranging from climate modeling to engineering design.
2 Background and Related Work
The emergence of neural operators has sparked significant advancements in scientific machine learning, particularly in the realm of solving partial differential equations (PDEs) and modeling complex physical systems. Building upon the foundational architectures mentioned in the introduction, recent developments have focused on enhancing the expressiveness and efficiency of these models. For instance, extensions of the Fourier Neural Operator (FNO) (Li et al., 2021) have explored factorized representations (Tran et al., 2021) to reduce computational complexity while maintaining performance. Similarly, wavelet-based approaches (Gupta et al., 2021; Tripura & Chakraborty, 2022) have been proposed to capture multi-scale features more effectively. Most existing neural operator architectures can be interpreted as conditioned neural fields (Xie et al., 2022) with specific choices of base fields and conditioning mechanisms, see Appendices B and C for a detailed discussion.
Concurrently, the success of vision transformers in computer vision tasks has inspired their adaptation to scientific machine learning problems. The self-attention mechanism central to transformers offers a powerful tool for capturing long-range dependencies in spatial and temporal data, a crucial aspect in many PDE-governed systems. However, adapting vision transformers to the continuous nature of physical problems poses unique challenges, particularly in handling irregular domains and ensuring output consistency across different resolutions.
In the specific context of operator learning, several transformer-based approaches have emerged to address these challenges. OFormer (Li et al., 2022) introduced a novel way to embed continuous input functions and output queries into a transformer architecture. This approach demonstrated the potential of transformers in handling the inherent continuity of physical problems. GNOT (Hao et al., 2023) further extended this idea by incorporating graph structures, allowing for more flexible representation of input functions and improved handling of irregular domains. A significant advancement in this direction came with DPOT (Hao et al., 2024), which introduced a denoising pre-training strategy. This approach enables the model to learn from diverse PDE datasets, significantly enhancing its generalization capabilities. However, these methods often require extensive computational resources for training and may struggle with maintaining consistency when evaluating at arbitrary resolutions.
Despite these advancements, several key challenges remain in the field of operator learning. These include efficiently handling high-dimensional data while capturing long-range dependencies, ensuring output consistency and accuracy across different resolutions, and balancing model expressiveness with computational efficiency. Addressing even a subset of these challenges requires novel architectural designs that can leverage the strengths of both neural operators and vision transformers while overcoming their respective limitations. This sets the stage for our proposed Continuous Vision Transformer (CViT), which aims to tackle specific challenges related to resolution independence and efficient capturing of multi-scale features in regular grid-based problems, combining the continuous nature of neural operators with the powerful representation capabilities of vision transformers.
@@ 3 Continuous Vision Transformer (CViT)
Figure 1: Continuous Vision Transformer (CViT) Architecture: CViT consists of the following components: (1) Spatio-temporal patch embeddings to extract localized features. (2) A temporal aggregation module based on the Perceiver architecture, which captures temporal correlations to compresses tokens along the time axis. (3) A Transformer encoder that captures multi-scale spatial dependencies via self-attention layers. (4) A novel grid-based positional encoding scheme for query coordinates, allowing for flexible output representation and interpolation. (5) A cross-attention decoder that integrates information from the input function with query coordinates.
While neural operators have shown promising results in learning mappings between function spaces, their architectural design has not fully leveraged recent advancements in deep learning. Popular architectures like FNO (Li et al., 2021), DeepONet (Lu et al., 2021), and NoMaD (Seidman et al., 2022) (see Appendix C) primarily rely on simple building blocks such as Fourier layers or fully connected networks, which may not effectively capture complex spatial dependencies and multiscale features present in many physical systems.
Motivated by the success of vision transformers in computer vision tasks, we introduce the Continuous Vision Transformer (CViT), a novel neural operator architecture that combines the strengths of vision transformers and continuous function representations. CViT aims to learn more powerful and flexible representations of time-dependent physical systems by leveraging a vision transformer encoder to capture complex spatial dependencies in input functions, a trainable grid-based positional encoding for flexible representation of query coordinates, and a cross-attention mechanism to integrate input function information with query coordinates.
Taken together, the overall design of CViT’s architecture draws inspiration from the analysis of existing neural operators presented in Appendix C, while leveraging the power of vision transformers to enhance expressiveness.
3.1 Architecture Description
The CViT architecture consists of three main components: a vision transformer encoder, a grid-based positional encoding, and a cross-attention mechanism.
Patch Embeddings.
The vision transformer encoder takes as input a gridded representation of the input function $u$, yielding a 2+1 dimensional spatio-temporal data tensor $\mathbf{u}\in{\mathbb{R}}^{T\times H\times W\times D}$ with $D$ channels. We patchify our inputs into 3D tokens $\mathbf{u}_{p}\in{\mathbb{R}}^{T\times\frac{H}{P}\times\frac{W}{P}\times C}$ by tokenizing each 2D spatial frame independently, following the process used in standard Vision Transformers (Dosovitskiy et al., 2020). We then add trainable 1D temporal and 2D spatial positional embeddings to each token:
Temporal Aggregation.
To reduce computational cost, we use a temporal aggregation layer based on the Perceiver architecture (Jaegle et al., 2021; Alayrac et al., 2022). Specifically, we learn a predefined number of latent input queries $\mathbf{z}\in{\mathbb{R}}^{1\times C}$ by applying a cross-attention module that maps the flattened visual features $\mathbf{u}_{f}\in{\mathbb{R}}^{(\frac{H}{P}\times\frac{W}{P})\times T\times C}$ to a latent tensor $\mathbf{z}_{agg}\in{\mathbb{R}}^{(\frac{H}{P}\times\frac{W}{P})\times 1\times C}$ as
$\mathbf{z}_{agg}$ $=\mathbf{z}^{\prime}+\operatorname{MLP}\left(\operatorname{LN}\left(\mathbf{z}^{\prime}\right)\right).$
Here, $\mathbf{z}$ is initialized by a unit Gaussian distribution and $\hat{\mathbf{z}}\in{\mathbb{R}}^{(\frac{H}{P}\times\frac{W}{P})\times T\times C}$ is obtained by tiling the latent query $\mathbf{z}$ $\left(\frac{H}{P}\times\frac{W}{P}\right)$ times. Besides compressing tokens over the time axis, this module also enables the model to handle inputs with a variable number of time steps.
Transformer Encoder.
We then process the aggregated tokens $\mathbf{z}_{agg}$ using a sequence of $L$ pre-norm Transformer blocks (Vaswani et al., 2017; Xiong et al., 2020),
$\mathbf{z}_{0}$ $=\operatorname{LN}(\mathbf{z}_{agg}),$
$\mathbf{z}_{\ell}^{\prime}$ $=\operatorname{MSA}\left(\operatorname{LN}\left(\mathbf{z}_{\ell-1}\right)\right)+\mathbf{z}_{\ell-1},$ $\ell=1\ldots L,$
Cross-Attention Decoder.
To enable continuous evaluation of outputs, we design a novel and efficient coordinate embedding to capture the fine-scale features of the target functions. Specifically, we create a uniform grid $\{\mathbf{y}_{ij}\}\subset[0,1]^{2}$, for $i=1,\dots N_{x}$ and $j=1,\dots N_{y}$, along with associated trainable latent grid features $\mathbf{x}\in{\mathbb{R}}^{N_{x}\times N_{y}\times C}$. For a query point $y\in{\mathbb{R}}^{2}$, we compute a Nadaraya-Watson interpolant over grid latent features:
$\mathbf{x}^{\prime}=\sum_{i=1}^{N_{x}}\sum_{j=1}^{N_{y}}w_{ij}\mathbf{x}_{ij},\quad w_{ij}=\frac{\exp\left(-\beta\|y-\mathbf{y}_{ij}\|^{2}\right)}{\sum_{ij}\exp\left(-\beta\|y-\mathbf{y}_{ij}\|^{2}\right)}.$ (1)
Here $\beta>0$ is a hyperparameter that determines the locality of the interpolated features. It is important for determining the smoothness of the interpolation function. Specifically, larger values of $\beta$ yield more localized weight distributions $w_{ij}$, resulting in a higher-frequency interpolant that captures finer-scale variations. Conversely, smaller $\beta$ values produce a smoother interpolant by averaging over a broader neighborhood of points.
The interpolated grid feature $\mathbf{x}_{0}=\mathbf{x}^{\prime}\in{\mathbb{R}}^{1\times C}$ is used as query input to a transformer decoder. This decoder uses the output of the vision transformer encoder $\mathbf{z}_{L}$ as keys and values in a cross-attention mechanism: